# TESLA-X: An effective method to search for sub-threshold lensed gravitational waves with a targeted population model

### v2

@@ Abstract
I Introduction
@@
Gravitational lensing refers to the phenomenon in which properties of propagating waves are influenced by the distortion of spacetime caused by massive intervening objects. The effects of gravitational lensing depends on the mass properties of the lens Ohanian (1974); Thorne (1982); Deguchi and Watson (1986); Wang et al. (1996); Nakamura (1998); Takahashi and Nakamura (2003). For massive lenses significantly larger than the gravitational-wave wavelength, we can assume geometric optics, which results in repeated signals with relative (de-)magnifications, time delays, and phase shifts Wang et al. (1996); Dai and Venumadhav (2017); Ezquiaga et al. (2021); Ng et al. (2018); Li et al. (2018); Oguri (2018); Smith et al. (2018, 2017, 2019); Robertson et al. (2020); Ryczanowski et al. (2020); Ezquiaga et al. (2023). Lenses with sizes comparable to gravitational-wave wavelengths can produce beating patterns through frequency-dependent wave optics effects, formally known as microlensing Deguchi and Watson (1986); Nakamura (1998); Takahashi and Nakamura (2003); Cao et al. (2014, 2014); Lai et al. (2018); Christian et al. (2018); Dai et al. (2018); Diego et al. (2019); Çalışkan et al. (2023a); Cheung et al. (2021); Mishra et al. (2021); Meena et al. (2022); Cremonese et al. (2021); Çalışkan et al. (2023b).
@@
The search for gravitationally-lensed gravitational waves is driven by various scientific objectives, such as testing general relativity Ezquiaga and Zumalacárregui (2020); Baker and Trodden (2017); Collett and Bacon (2017); Goyal et al. (2023); Ezquiaga et al. (2021), improving source localization accuracy Hannuksela et al. (2020); Yu et al. (2020), characterizing lens models Lai et al. (2018); Diego et al. (2019); Oguri and Takahashi (2020); Liu et al. (2023), and performing precision cosmology Sereno et al. (2011); Liao et al. (2017); Cao et al. (2019); Hannuksela et al. (2020) and statistical cosmology Xu et al. (2022).
@@
The LIGO-Virgo-KAGRA (LVK) collaboration has conducted various analyses searching for lensing signatures in the event catalogs of LVK’s observing runs O1-O2 Hannuksela et al. (2019) and O3 Abbott et al. (2021a, 2023), but no convincing evidence of lensing signatures has been found. Additionally, other analyses have concluded that there are no confident lensing signatures in the detected gravitational-wave events Li et al. (2023); McIsaac et al. (2020); Dai et al. (2020); Liu et al. (2021); Ezquiaga et al. (2023).
@@
Strongly lensed gravitational waves will likely originate from sources living in the higher redshift universe. Without being lensed, these signals will only be barely detectable or undetectable. Under strong lensing, copies of gravitational waves from these high redshift sources will be produced; some may be magnified and are detected as super-threshold gravitational waves, but some may be potentially de-magnified with a reduced signal-to-noise (SNR) ratio below the detection threshold. These signals, formally known as “sub-threshold” signals, are of particular interest, as a significant proportion of strongly-lensed gravitational-wave events falls into this category Wierda et al. (2021). Retrieving these possible sub-threshold lensed counterparts is a crucial step in boosting confidence for the first detection of strongly-lensed gravitational waves.
@@
Two targeted search methods Li et al. (2023); McIsaac et al. (2020) have been developed and employed to recover potential sub-threshold lensed counterparts to known super-threshold gravitational waves. The GstLAL-based TargetEd Subthreshold Lensing seArch (TESLA) described in Li et al. (2023) strategically reduces the search parameter space to look for possible sub-threshold lensed counterparts. However, as we will explain in this paper, the traditional TESLA method can potentially lose signals due to an excessive loss in signal-to-noise (SNR) ratio due to how the targeted template bank is constructed.
@@
Additionally, TESLA’s efficiency in finding potential sub-threshold lensed gravitational waves can be improved if a targeted population model built based on information from the target superthreshold event is used. A proper targeted population model in the ranking statistic can further reduce the influence of the noise background. Detailed information on the TESLA method can be found in Li et al. (2023).
@@
This paper is organized as follows: Section II provides a brief review of strong lensing of gravitational waves. Section III outlines the workflow of the traditional TESLA method. In Section IV, we give an overview of how the population prior term in the likelihood ratio calculations for triggers from the GstLAL search pipeline is calculated, and detail the TESLA-X methodology. Section V presents a simulation campaign to evaluate the TESLA-X method. Finally, in Section VI, we summarize our findings and discuss potential future work to further enhance the search sensitivity of the TESLA-X pipeline.
II Strong lensing of gravitational waves
@@
Under the strong lensing hypothesis, strong gravitational lenses can produce multiple copies of transient gravitational waves from the same source. These repeated signals have identical waveforms (i.e. their underlying intrinsic parameters $\vec{\theta}$, e.g. component masses and spins, are the same), except that (I) their amplitudes may differ by some amplitude scaling factor $\sqrt{\mu_{j}}$ (Note that the amplitude scaling factor is achromatic under the geometric optics assumption, and the amplification factor can be $>1$ (magnified) or $\leq 1$ (de-magnified).), (II) they are separated by a relative arrival time delay $\Delta t_{j}$ with respect to the not-lensed signal’s coalescence time, and (III) they have an additional Morse phase factor $\Delta\phi_{j}$ applied to depending on the type of lensed signals (Note: There are three types of lensed signals, type I, II and III. They correspond to the minimum, saddle point, and maximum time-delay solution to the lens equation, respectively.), given by Ezquiaga and Zumalacárregui (2020); Dai et al. (2020)
$\Delta\phi_{j}=-\frac{n_{j}\pi}{2},\,\,n_{j}=\begin{cases}0,\text{Type I lensed signals}\\ 1,\text{Type II lensed signals}\\ 2,\text{Type III lensed signals}\end{cases}.$ (1)
@@
Since the travelling paths of these repeated signals have been altered, these signals will appear to be originating from sources at different sky locations. However, the deviation in sky location due to strong lensing is in order of arc-seconds. This is negligible compared to the uncertainty in sky localization for gravitational waves, which is in the order of degrees Abbott et al. (2018). Therefore, for the rest of this paper we will assume multiple strongly-lensed gravitational-wave signals from the same source to have the same sky location.
@@
III The search for sub-threshold strongly-lensed gravitational waves: Objectives and current approach
@@
III.1 Current effort to search for strongly-lensed sub-threshold lensed counterparts to super-threshold gravitational waves
@@
When searching for gravitational waves in general, signals with sufficiently high amplitudes that can be identified as gravitational waves from the noisy data are known as super-threshold. In contrast, weaker gravitational waves not distinguishable from noise are known as sub-threshold. As discussed in Section II, strong lensing can produce multiple copies of gravitational waves coming from the same source with different amplification factors $\sqrt{\mu_{j}}$. $\sqrt{\mu_{j}}$ can take on values smaller than $1$, i.e. the repeated gravitational-wave signals are de-magnified compared to the not-lensed gravitational-wave signal and may become sub-threshold. It is then natural to ask the question: If we assume that one of the identified super-threshold gravitational waves detected so far is strongly lensed, can we find its sub-threshold strongly lensed counterparts, if they exist?
@@
There have been ongoing efforts to employ matched-filtering based search pipelines with modifications Li et al. (2023); McIsaac et al. (2020) to search for strongly-lensed sub-threshold lensed counterparts to identified super-threshold gravitational waves from LVK’s first three observing runs (O1, O2 and O3) Hannuksela et al. (2019); Abbott et al. (2021a, 2023). In this paper, we focus on the TargetEd subthreshold Lensing seArch (TESLA) method Li et al. (2023), which is built based on the GstLAL search pipeline Messick et al. (2017); Sachdev et al. (2019); Hanna et al. (2020); Allen et al. (2012); Cannon et al. (2021); Godwin et al. (2020); Chan et al. (2020); Mukherjee et al. (2021); Messick et al. (2020). The following sub-section gives a brief introduction to the traditional TESLA method.
III.2 The traditional TESLA search method
@@
Traditional TESLA is built based on the matched-filtering-based pipeline GstLAL. A detailed introduction to the GstLAL search pipeline can be found in Li et al. (2023); Messick et al. (2017); Cannon et al. (2021). In a general search for gravitational waves from compact binary coalescences, the goal is to detect signals that span the entire parameter space to which the detector network’s frequency bandwidth is sensitive. Hence, a large template bank comprising over a million template waveforms is constructed and used. A large number of candidates (also known as triggers) is generated, regardless of whether they are noise or signals. The more templates we use in a search, the larger the trials factor, and hence, the larger the noise background will become.
@@
Assuming a super-threshold gravitational wave is strongly lensed, we want to search for its possible sub-threshold lensed counterparts. In section II, we explained that strongly-lensed gravitational waves from the same source should have identical intrinsic parameters (e.g. component masses and spins), and hence identical waveforms. For each super-threshold event (target), we obtain the posterior probability distribution that gives the best estimates of the source parameters of the target using the Bayesian parameter estimation analyses described in Veitch et al. (2015); Ashton et al. (2019); Romero-Shaw et al. (2020); Ashton and Talbot (2021). We can then narrow down the parameter search space for possible strongly-lensed sub-threshold counterparts to regions consistent with the target’s posterior parameter space. However, as explained in Li et al. (2023), the posterior parameter space of the target itself is insufficient because of noise fluctuations in the data. Non-Gaussianity in the data can cause false signals (especially those that have weaker amplitudes and are hence sub-threshold), to be registered by templates that have very different parameters than those of the posterior samples for the target.
TESLA accounts for both the signal sub-space and noise fluctuations in the data when searching for possible strongly-lensed sub-threshold counterparts to super-threshold gravitational waves. Here, we outline the steps for running a TESLA analysis. For each super-threshold gravitational event (target), we start with the posterior samples ordered in decreasing order of log-likelihood from the parameter estimation for the target. Since the amplitudes of possible strongly-lensed sub-threshold gravitational waves are lower, they will be registered as triggers in the search pipeline with lower signal-to-noise ratio $\rho$ (SNR). For each posterior sample, we generate $1$ injection with the exact same parameters (including luminosity distance to the source $D_{L}$ and sky location, i.e. right ascension $\alpha$ and declination $\delta$) as the sample, and $9$ extra injections with $D_{L}$ increased (and hence optimal SNR $\rho_{\mathrm{opt}}$ decreased) to mimic the de-magnifying effect of strong lensing on the amplitudes of possible sub-threshold lensed counterparts, according to the inverse proportionality of $D_{L}$ and the optimal SNR $\rho_{\mathrm{opt}}$:
@@
We require the latter $9$ injections to have detector SNR $\rho_{\mathrm{det}}\geq 4$ in at least one detector as a constraint from the GstLAL search pipeline (Note: We do not lower the single detector SNR threshold in the GstLAL search pipeline since this will lead to an exponential increase in the number of both signal and noise triggers. As future work, we will explore the possibility of further lowering the single detector SNR threshold.). The set of lensed injections represents the possible sub-threshold lensed counterparts to the target we are searching for. They are then injected into real data, and we perform an injection campaign using GstLAL with a full template bank to recover them. After the injection campaign, templates that can find these injections (Note: An injection is “found” if the FAR of the associated trigger is $\leq$ $1$ in $30$ days.) are kept to construct a targeted template bank to search for possible strongly-lensed sub-threshold counterparts to the target. The end product of TESLA is a ranked list of possible lensed counterparts to the target in increasing order of FARs. Readers are reminded that the FARs (or other ranking statistics) reported in the search’s ranked list do not indicate how likely the signal candidate is a lensed counterpart to the target, but only a priority label for follow-up analysis.
For a detailed explanation of the TESLA method, please refer to Li et al. (2023).
IV The upgraded TESLA-X search method
IV.1 The missing pieces in the traditional TESLA method: Non-optimal template bank and population model
We have shown in Li et al. (2023) that the traditional TESLA method succeeds in improving the search sensitivity to possible strongly-lensed sub-threshold gravitational waves. However, the whole picture still needs to include two puzzle pieces.
First, in the original full template bank, the parameter ranges are informed by the masses and spins of detectable sources (BNS, NSBH, BBH), and the templates are distributed with a density that allows a maximum loss of $3\%$ SNR when a gravitational-wave signal is recovered (compared to the optimal SNR (Note: The optimal SNR is the SNR when the signal is cross-correlated with a template with the exact same parameters as the signal) Mukherjee et al. (2021). However, when a reduced template bank is generated using traditional TESLA, we discard templates that did not recover lensed injections for the target super-threshold gravitational waves. This breaks the original optimality of the full template bank, caused by the finite number of lensed injections that are used to generate the reduced template bank, and strongly-lensed sub-threshold signals whose optimal detector SNR is near threshold (i.e. around $\rho_{\mathrm{opt}}\approx 4.12$) will not be detected. Therefore, it is vital that the optimality of the reduced template bank must be conserved to maximize our search sensitivity towards possible strongly-lensed sub-threshold gravitational waves.

### v5

@@ Abstract
keywords:
gravitational lensing – gravitational waves – black hole mergers
1 Introduction
@@
Gravitational lensing refers to the phenomenon in which properties of propagating waves are influenced by the distortion of spacetime caused by massive intervening objects. The effects of gravitational lensing depends on the mass properties of the lens (Ohanian, 1974; Thorne, 1982; Deguchi & Watson, 1986; Wang et al., 1996; Nakamura, 1998; Takahashi & Nakamura, 2003). For massive lenses significantly larger than the gravitational-wave wavelength, we can assume geometric optics, which results in repeated signals with relative (de-)magnifications, time delays, and phase shifts (Wang et al., 1996; Dai & Venumadhav, 2017; Smith et al., 2017; Smith et al., 2018; Smith et al., 2019; Li et al., 2018; Ng et al., 2018; Oguri, 2018; Robertson et al., 2020; Ryczanowski et al., 2020; Ezquiaga et al., 2021; Janquart et al., 2021; Ezquiaga et al., 2023; Lo & Magana Hernandez, 2023; Barsode et al., 2025; Chan et al., 2025b). Lenses with sizes comparable to gravitational-wave wavelengths can produce beating patterns through frequency-dependent wave optics effects, formally known as microlensing (Deguchi & Watson, 1986; Nakamura, 1998; Takahashi & Nakamura, 2003; Cao et al., 2014; Christian et al., 2018; Dai et al., 2018; Lai et al., 2018; Diego et al., 2019; Cheung et al., 2021; Cremonese et al., 2021; Mishra et al., 2021; Meena et al., 2022; Çalışkan et al., 2023a, b; Chan et al., 2025a).
@@
The search for gravitationally-lensed gravitational waves is driven by various scientific objectives, such as testing general relativity (Baker & Trodden, 2017; Collett & Bacon, 2017; Ezquiaga & Zumalacárregui, 2020; Ezquiaga et al., 2021; Goyal et al., 2023; Narayan et al., 2024), improving source localization accuracy (Hannuksela et al., 2020; Yu et al., 2020; Vujeva et al., 2025b, a), characterizing lens models (Lai et al., 2018; Diego et al., 2019; Oguri & Takahashi, 2020; Liu et al., 2023), and performing precision cosmology (Sereno et al., 2011; Liao et al., 2017; Cao et al., 2019; Hannuksela et al., 2020; Farah et al., 2025) and statistical cosmology (Xu et al., 2022).
@@
The LIGO-Virgo-KAGRA (LVK) collaboration has conducted various analyses searching for lensing signatures in the event catalogs of LVK’s observing runs O1-O2 (Hannuksela et al., 2019) and O3 (Abbott et al., 2021b, 2024), but no convincing evidence of lensing signatures has been found. Additionally, other analyses have concluded that there are no confident lensing signatures in the detected gravitational-wave events (Li et al., 2023; McIsaac et al., 2020; Dai et al., 2020; Liu et al., 2021; Ezquiaga et al., 2023). None the less, Wierda et al. (2021) predicted that the current LVK detector network at its design sensitivity can reach a lensing detection rate of $1.9$ events per year. Furthermore, Wang et al. (2021) projected that in the era of the Cosmic Explorer (CE), the annual rate of strongly lensed gravitational-wave detections could reach as high as $184.7$.
@@
Strongly lensed gravitational waves will likely originate from sources living in the higher redshift universe. Without being lensed, these signals will only be barely detectable or undetectable. Under strong lensing, copies of gravitational waves from these high redshift sources will be produced; some may be magnified and are detected as super-threshold gravitational waves, but some may be potentially de-magnified with a reduced signal-to-noise (SNR) ratio below the detection threshold. These signals, formally known as “sub-threshold” signals, are of particular interest, as a significant proportion of strongly-lensed gravitational-wave events falls into this category (Wierda et al., 2021). Retrieving these possible sub-threshold lensed counterparts is a crucial step in boosting confidence for the first detection of strongly-lensed gravitational waves.
@@
Two targeted search methods (Li et al., 2023; McIsaac et al., 2020) have been developed and employed to recover potential sub-threshold lensed counterparts to known super-threshold gravitational waves. The GstLAL-based TargetEd Subthreshold Lensing seArch (TESLA) described in Li et al. (2023) strategically reduces the search parameter space to look for possible sub-threshold lensed counterparts. However, as we will explain in this paper, the traditional TESLA method can potentially lose signals due to an excessive loss in signal-to-noise (SNR) ratio due to how the targeted template bank is constructed.
@@
Additionally, TESLA’s efficiency in finding potential sub-threshold lensed gravitational waves can be improved if a targeted population model built based on information from the target superthreshold event is used. A proper targeted population model in the ranking statistic can further reduce the influence of the noise background. Detailed information on the TESLA method can be found in (Li et al., 2023).
@@
This paper is organized as follows: Section 2 provides a brief review of strong lensing of gravitational waves. Section 3 outlines the workflow of the traditional TESLA method. In Section 4, we give an overview of how the population prior term in the likelihood ratio calculations for triggers from the GstLAL search pipeline is calculated, and detail the TESLA-X methodology. Section 5 presents a simulation campaign to evaluate the TESLA-X method. Finally, in Section 6, we summarize our findings and discuss potential future work to further enhance the search sensitivity of the TESLA-X pipeline.
2 Strong lensing of gravitational waves
@@
Under the strong lensing hypothesis, strong gravitational lenses can produce multiple copies of transient gravitational waves from the same source. These repeated signals have identical waveforms (i.e. their underlying intrinsic parameters $\vec{\theta}$, e.g. component masses and spins, are the same), except that (I) their amplitudes may differ by some amplitude scaling factor $\sqrt{\mu_{j}}$ (Note that the amplitude scaling factor is achromatic under the geometric optics assumption, and the amplification factor can be >1>1 (magnified) or $\leq 1$ (de-magnified).), (II) they are separated by a relative arrival time delay $\Delta t_{j}$ with respect to the not-lensed signal’s coalescence time, and (III) they have an additional Morse phase factor $\Delta\phi_{j}$ applied to depending on the type of lensed signals (Note: There are three types of lensed signals, type I, II and III. They correspond to the minimum, saddle point, and maximum time-delay solution to the lens equation, respectively.), given by Ezquiaga & Zumalacárregui (2020); Dai et al. (2020)
[…]
@@
Since the travelling paths of these repeated signals have been altered, these signals will appear to be originating from sources at different sky locations. However, the deviation in sky location due to strong lensing is in order of arc-seconds. This is negligible compared to the uncertainty in sky localization for gravitational waves, which is in the order of degrees (Abbott et al., 2018). Therefore, for the rest of this paper we will assume multiple strongly-lensed gravitational-wave signals from the same source to have the same sky location.
@@
3 The search for sub-threshold strongly-lensed gravitational waves: Objectives and current approach
@@
3.1 Current effort to search for strongly-lensed sub-threshold lensed counterparts to super-threshold gravitational waves
@@
When searching for gravitational waves in general, signals with sufficiently high amplitudes that can be identified as gravitational waves from the noisy data are known as super-threshold. In contrast, weaker gravitational waves not distinguishable from noise are known as sub-threshold. As discussed in Section 2, strong lensing can produce multiple copies of gravitational waves coming from the same source with different amplification factors $\sqrt{\mu_{j}}$. $\sqrt{\mu_{j}}$ can take on values smaller than $1$, i.e. the repeated gravitational-wave signals are de-magnified compared to the not-lensed gravitational-wave signal and may become sub-threshold. It is then natural to ask the question: If we assume that one of the identified super-threshold gravitational waves detected so far is strongly lensed, can we find its sub-threshold strongly lensed counterparts, if they exist?
@@
There have been ongoing efforts to employ matched-filtering based search pipelines with modifications (Li et al., 2023; McIsaac et al., 2020) to search for strongly-lensed sub-threshold lensed counterparts to identified super-threshold gravitational waves from LVK’s first three observing runs (O1, O2 and O3) (Hannuksela et al., 2019; Abbott et al., 2021b, 2024). In this paper, we focus on the TargetEd subthreshold Lensing seArch (TESLA) method (Li et al., 2023), which is built based on the GstLAL search pipeline (Messick et al., 2017; Sachdev et al., 2019; Hanna et al., 2020; Allen et al., 2012; Cannon et al., 2021; Godwin et al., 2020; Chan et al., 2020; Mukherjee et al., 2021; Messick et al., 2020). The following sub-section gives a brief introduction to the traditional TESLA method.
3.2 The traditional TESLA search method
@@
Traditional TESLA is built based on the matched-filtering-based pipeline GstLAL. A detailed introduction to the GstLAL search pipeline can be found in (Li et al., 2023; Messick et al., 2017; Cannon et al., 2021). In a general search for gravitational waves from compact binary coalescences, the goal is to detect signals that span the entire parameter space to which the detector network’s frequency bandwidth is sensitive. Hence, a large template bank comprising over a million template waveforms is constructed and used. A large number of candidates (also known as triggers) is generated, regardless of whether they are noise or signals. The more templates we use in a search, the larger the trials factor, and hence, the larger the noise background will become.
@@
Assuming a super-threshold gravitational wave is strongly lensed, we want to search for its possible sub-threshold lensed counterparts. In section 2, we explained that strongly-lensed gravitational waves from the same source should have identical intrinsic parameters (e.g. component masses and spins), and hence identical waveforms. For each super-threshold event (target), we obtain the posterior probability distribution that gives the best estimates of the source parameters of the target using the Bayesian parameter estimation analyses described in (Veitch et al., 2015; Ashton et al., 2019; Romero-Shaw et al., 2020; Ashton & Talbot, 2021). We can then narrow down the parameter search space for possible strongly-lensed sub-threshold counterparts to regions consistent with the target’s posterior parameter space. However, as explained in (Li et al., 2023), the posterior parameter space of the target itself is insufficient because of noise fluctuations in the data. Non-Gaussianity in the data can cause false signals (especially those that have weaker amplitudes and are hence sub-threshold), to be registered by templates that have very different parameters than those of the posterior samples for the target.
TESLA accounts for both the signal sub-space and noise fluctuations in the data when searching for possible strongly-lensed sub-threshold counterparts to super-threshold gravitational waves. Here, we outline the steps for running a TESLA analysis. For each super-threshold gravitational event (target), we start with the posterior samples ordered in decreasing order of log-likelihood from the parameter estimation for the target. Since the amplitudes of possible strongly-lensed sub-threshold gravitational waves are lower, they will be registered as triggers in the search pipeline with lower signal-to-noise ratio $\rho$ (SNR). For each posterior sample, we generate 11 injection with the exact same parameters (including luminosity distance to the source $D_{L}$ and sky location, i.e. right ascension $\alpha$ and declination $\delta$) as the sample, and 99 extra injections with $D_{L}$ increased (and hence optimal SNR $\rho_{\mathrm{opt}}$ decreased) to mimic the de-magnifying effect of strong lensing on the amplitudes of possible sub-threshold lensed counterparts, according to the inverse proportionality of $D_{L}$ and the optimal SNR $\rho_{\mathrm{opt}}$:
@@
We require the latter $9$ injections to have detector SNR $\rho_{\mathrm{det}}\geq 4$ in at least one detector as a constraint from the GstLAL search pipeline (Note: We do not lower the single detector SNR threshold in the GstLAL search pipeline since this will lead to an exponential increase in the number of both signal and noise triggers. As future work, we will explore the possibility of further lowering the single detector SNR threshold.). The set of lensed injections represents the possible sub-threshold lensed counterparts to the target we are searching for. They are then injected into real data, and we perform an injection campaign using GstLAL with a full template bank to recover them. After the injection campaign, templates that can find these injections (Note: An injection is “found” if the FAR of the associated trigger is $\leq$ 11 in $30$ days.) are kept to construct a targeted template bank to search for possible strongly-lensed sub-threshold counterparts to the target. The end product of TESLA is a ranked list of possible lensed counterparts to the target in increasing order of FARs. Readers are reminded that the FARs (or other ranking statistics) reported in the search’s ranked list do not indicate how likely the signal candidate is a lensed counterpart to the target, but only a priority label for follow-up analysis.
For a detailed explanation of the TESLA method, please refer to (Li et al., 2023).
4 The upgraded TESLA-X search method
4.1 The missing pieces in the traditional TESLA method: Non-optimal template bank and population model
We have shown in (Li et al., 2023) that the traditional TESLA method succeeds in improving the search sensitivity to possible strongly-lensed sub-threshold gravitational waves. However, the whole picture still needs to include two puzzle pieces.
First, in the original full template bank, the parameter ranges are informed by the masses and spins of detectable sources (BNS, NSBH, BBH), and the templates are distributed with a density that allows a maximum loss of $3\%$ SNR when a gravitational-wave signal is recovered (compared to the optimal SNR (Note: The optimal SNR is the SNR when the signal is cross-correlated with a template with the exact same parameters as the signal) (Mukherjee et al., 2021). However, when a reduced template bank is generated using traditional TESLA, we discard templates that did not recover lensed injections for the target super-threshold gravitational waves. This breaks the original optimality of the full template bank, caused by the finite number of lensed injections that are used to generate the reduced template bank, and strongly-lensed sub-threshold signals whose optimal detector SNR is near threshold (i.e. around $\rho_{\mathrm{opt}}\approx 4.12$) will not be detected. Therefore, it is vital that the optimality of the reduced template bank must be conserved to maximize our search sensitivity towards possible strongly-lensed sub-threshold gravitational waves.